OpenStudy (anonymous):

If y = 5.7 when x = –3.99, what is y when x = –0.49? Answer 0.34 –0.34 –0.7 0.7

OpenStudy (anonymous):

y1 5.7 y2 -- = ------ = ------ x1 -3.99 x2 so 5.7 y2 ------ = ------ -3.99 -0.49 this gives 5.7 * -0.49 ---------- = y2 -3.99 y2 = 0.7 so the answer is 0.7